Verify that the total differential equation(1+yz)dx+x(z-x)dy-(1+xy)dz=0 is integrable and hence find its integral.


1
Expert's answer
2021-02-01T07:23:30-0500

Necessary and sufficient condition for the total differential equation "Pdx+Qdy+Rdz=0"


to be integrable is "P(\u2202R\/\u2202y-\u2202Q\/\u2202z)-Q(\u2202R\/\u2202x-\u2202P\/\u2202z)+\n+R(\u2202Q\/\u2202x-\u2202P\/\u2202y)=0"

"P=1+yz"

"\u2202P\/\u2202z=y"

"\u2202P\/\u2202y=z"

"Q=x(z-x)"

"\u2202Q\/\u2202x=z-2x"

"\u2202Q\/\u2202z=x"

"R=-(1+xy)"

"\u2202R\/\u2202x=-y"

"\u2202R\/\u2202y=-x"


According to the given equation

"(1+yz)(-x-x)-x(z-x)(-y-y)+(-1-xy)(z-2x-z)=\\
=-2x-2xyz+2xyz-2x^2 y+2x+2x^2 y=0"


So it is integrable.

 

Let "y=const"

"(1+yz)dx-(1+xy)dz=0"

Integrate

"x+xyz-z-xyz=f(y)"

"x-z=f(y)"

Use the equation

"(1+yz)dx+f' (y)dy-(1+xy)dz=0"

"f' (y)=x(z-x)"

"f' (y)=-f(y)x"

"ln\u2061\u3016f(y)\u3017=-xy+C"

"f(y)=exp(-xy+C)"

Answer: "x-z=exp(-xy+C)"